What is the magnitude and direction of the electric field at 0.25 m from a -5.6x10-6 C point charge? Choose the correct answer.


a. 1.18x1055 N/C toward the charge

b. 1.18x105 N/C away from the charge

c. 8.10x105 N/C toward the charge


1
Expert's answer
2021-03-21T11:27:20-0400

We can find the magnitude of the electric field as follows:


"E=\dfrac{k|q|}{r^2}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot|-5.6\cdot10^{-6}\ C|}{(0.25\ m)^2}=8.1\cdot10^5\ \dfrac{N}{C}."

The electric field directed toward the charge.

Answer:

c. 8.10x10N/C toward the charge.